xx IDFileADM

IDFileADM

Erik Schnetter <schnetter@aei.mpg.de>

Date

Abstract

This thorn makes it possible to use the file reader to read initial data for the ADM variables. It only makes it possible to use the file reader; it does not actually activate the file reader, which has to be done by hand. This thorn also checks whether the ADM variables are actually supposed to be read, so that it is impossible to actually leave something uninitialised. However, read errors are also not caught.

1 Parameters

11conformalxstorage 10Scope: shared from STATICCONFORMALx01KEYWORD

2 Interfaces

General Implements: idfileadm

Inherits: admbase staticconformal io

3 Schedule This section lists all the variables which are assigned storage by thorn EinsteinInitialData/IDFileADM. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK x idfileadmxparamcheck x ensure that all admbase initial data are either initialised or read from a file x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxINITIAL x idfileadmxinit x initialise the conformalxstate x

11x11Language:11c
11 x 11Type: 11function